During an episode of turbulence in an airplane, you feel 170 N heavier than usual. If your mass is 86 kg, what are the magnitude and direction of the airplane’s acceleration?

Ims b) 3.00 ms2, downward c) 1.98 ms2, upward C> d) 2.25 ms^, upward
Answer #1

Weight= 86*9.8= 842.8 N

additional weight= 170 N

total apparent weight= 842.8+ 170 =1012.8 N

using newton's second law

1012.8-842.8 = 86*a

170/86= a

a=1.98 m/s2, upward so option c